# Less is more - the dispatcher/ executor principle for multi-task Reinforcement Learning

### v1

@@ Abstract
Humans instinctively know how to neglect details when it comes to solve complex decision making problems in environments with unforeseeable variations. This abstraction process seems to be a vital property for most biological systems and helps to ’abstract away’ unnecessary details and boost generalisation. In this work we introduce the dispatcher/ executor principle for the design of multi-task Reinforcement Learning controllers. It suggests to partition the controller in two entities, one that understands the task (the dispatcher) and one that computes the controls for the specific device (the executor) - and to connect these two by a strongly regularizing communication channel. The core rationale behind this position paper is that changes in structure and design principles can improve generalisation properties and drastically enforce data-efficiency. It is in some sense a ’yes, and …’ response to the current trend of using large neural networks trained on vast amounts of data and bet on emerging generalisation properties. While we agree on the power of scaling - in the sense of Sutton’s ’bitter lesson’ - we will give some evidence, that considering structure and adding design principles can be a valuable and critical component in particular when data is not abundant and infinite, but is a precious resource. A video showing the results can be found at https://sites.google.com/view/dispatcher-executor.
@@ 1 Introduction
Reinforcement Learning (RL) has matured from single task settings to multi-task scenarios, where a single controller is asked to solve one of many tasks. Multi-task requirements commonly occur in robotics, where e.g. a robot arm is asked to manipulate a wide range of objects in different ways, or a locomotion platform should reach different goals. Multi-task requirements also occur in classical control applications, e.g. controlling fusion plasma to various shapes (Degrave et al., 2022) or set-point control problems in general (Hafner & Riedmiller, 2011). The typical approach for multi-task control is to use the ’classical’ controller structure, represented by a monolithic neural net, that is extended by a pathway to specify the requested task.
However, it requires different kinds of knowledge for different aspects of the control task: Task understanding, e.g. understanding which objects in a scene are affected or which type of manipulation is requested, requires a general understanding of the world, whereas computing a concrete control action requires specific knowledge about the device at hand.
In this paper we argue, that multi-task RL controllers can massively benefit from a bi-partite controller structure, that cleanly separates the overall architecture into one part that semantically understands the task (the ’dispatcher’) and a second part that executes the input-to-action policy to control the device (the ’executor’). A further crucial point is that the communication between dispatcher and executor should enforce compositionality and a substantial reduction of irrelevant information about the situation. This forced abstraction from the details allows the same executor to be applied in a broad diversity of original scenes and thus boosts generalisation - hence the title ’less is more’.
We propose the dispatcher/ executor (D/E) principle for multi-task control and show a concrete implementation in the domain of robotic manipulation and discuss its properties. We provide a set of empirical evaluations and show their significant benefits on a range of typical robotic manipulation tasks in both simulation and on the real robot. The dispatcher/ executor principle can even be applied in hindsight, i.e. if a controller for a certain concrete task has already been learned. We show this hindsight transfer to a D/E structure on the example of a real robot policy that is transferred from stacking ’red on blue’ objects to a general object stacking routine.
@@
  • introduces the dispatcher/ executor (D/E) principle as a design principle for multi-task RL controllers
  • makes a concrete proposal for such a dispatcher/ executor architecture in the area of robotic manipulation – both as a proof of principle and as a concrete practical implementation proposal
  • provides empirical evaluations in simulation and on a real robot demonstrating huge gains in data-efficiency in multi-task RL settings
@@ 2 State of the Art
The core claim of Rich Sutton’s insightful ’Bitter Lesson’ paper (Sutton, 2019) is simple and compelling: ’general methods that leverage computation are ultimately the most effective’. This has been demonstrated several times in the history of AI, e.g. in game playing (Silver et al., 2016), computer vision (Krizhevsky et al., 2017), or more recently, in large language models (Vaswani et al., 2017): huge amounts of data trained in huge neural networks with huge amounts of compute lead to unprecedented and surprising generalisation (OpenAI, 2023), while the underlying training principle is general and relatively simple.
@@
Given the big success in language, it is evident to investigate, if the same principles of scaling up to large amounts of data and the use of large networks show the same generalisation effects in control. The GATO (Reed et al., 2022) approach demonstrated that a reasonably large transformer style neural network is able to show good control performance with a single network on a variety of heterogeneous tasks from Atari to robotics. RoboCat (Bousmalis et al., 2023) used a related approach focusing on robotic control with the final goal to train a versatile model that can transfer to arbitrary new tasks. RT-1 (Brohan et al., 2022) and RT-2 (Zitkovich et al., 2023) start from a pre-trained language model and therefore have a built-in semantic understanding of tasks and the environment. All of the above models are trained on vast amounts of data using a transformer-style controller architecture with an additional input pathway to communicate the task. Since these methods are by default extremely data-hungry, data-augmentation is one path to generate even more data out of the data collected (Yu et al., 2023) and by this enforce generalisation within the distribution of the data generated.
@@
In Ahn et al. (2022) it is shown that a pre-trained language model and a set of pre-trained skills can be efficiently combined by learning to sequence these skills, based on the assumption that language model and the skills share some common understanding of the main task. While this gives some form of general task execution in known environments, a large amount of data has to be used to make sure the pre-trained skills will be able to execute on a sufficient enough set of variations of the environment.
@@
Another scheme that is very common for data efficient multi task approaches is to learn some form of task embedding that allows to generalise in the task space (e.g. Hausman et al. (2018); Nachum et al. (2018)). While some of these ideas can be combined with the D/E approach, these methods solely focus on the command part in the communication channel. Therefore the expected generalisation by abstracting the observation and the command, as we demonstrate here, is missing.
@@
While we agree that scaling is one way to achieve more and more general controllers, we at the same time believe that even large scale systems should be ’data-efficient’, i.e. should make the best use of the data available. The core thesis of this paper is that incorporating structure into the architecture and the communication channel can drastically enhance the generalisation capabilities ’by design’. In particular, enforcing the right representation within the controller seems to be an important enabler. Recent examples of this line of research can be found in Levine et al. (2016); Pinneri et al. (2023); Groth et al. (2021).
@@
The idea of a two (or more) level architectures for control has been explored in many different variations (e.g. Vezzani et al. (2022); Zhang & Chai (2021); Riedmiller et al. (2018); Hafner et al. (2022). The D/E principle intentionally defines the separation from the viewpoint of the type of knowledge required to run the components: a dispatcher component that understands the task through general world knowledge (that can be trained on general data like text, images or videos), and an executor component that specifically understands the device (trained through active interaction with the device). The restrictions on the communication channel enforce an abstract information exchange between these types of knowledge that boosts generalisation. We believe that in particular in this new era of large language models that are able to represent general world knowledge, this structural principle can play an important role to boost learning control on real robots, where data-efficiency with respect to device interaction is a key requirement.
@@ 3.3 Workflow and learning within the D/E architecture for manipulation
Learning in a D/E architecture is currently restricted to the executor. Learning is not different from the standard monolithical controller, apart from the fact, that the input is represented differently. The executor can therefore be trained in many different ways. In the following we investigate Reinforcement Learning from scratch as well as the distillation of a previously learned policy to a D/E architecture in a student/teacher setup. Another interesting direction, which is beyond the scope of this paper, is the investigation of learning multiple variants of representations in an off-line RL setting of pre-collected data (Lange et al., 2012; Riedmiller et al., 2021).
@@ 4 Empirical investigations of targeted properties
The dispatcher-executor architecture (see figure 1) uses a hardcoded dispatcher and a learned executor. The dispatcher gets the task specification (e.g. ’lift green object’ or ’stack red on blue object’) and then segments the according target objects - in this case using simple color segmentation (however, it is straightforward to replace this by more capable segmentation schemes, e.g. Oquab et al. (2023). In addition, it filters the whole scene using an edge operator. Then, the executor is called, providing the respectively filtered images as separate arguments. If not otherwise noted, ’D/E’ refers to the default setting of using a segmentation mask for the objects, and the edge operator for the overall scene (as described in 3.2).
For training both the standard architecture and the executor in the D/E case, a distributional variant of MPO Abdolmaleki et al. (2018), with a Mixture-of-Gaussian critic was used. The agent receives the images of three cameras placed around the robot’s workspace, as well as proprioceptive observations. Both architectures employ a three-block ResNet that computes an eight-dimensional embedding vector for each pixel input. These embeddings are concatenated with all proprioceptive observations and fed through a three layer MLP for computing the agent’s five-dimensional actions.
@@ 4.3 Multi-task 2: Lifting one out of three objects with multiple shapes (simulation)
One critical question with a reduced representation is, whether the carried information is still rich enough to cope for desired variances in the task setup. To look a bit deeper into this, in the following the objects are additionally varied in shape. This problem is considerably more complicated than grasping cubes, forcing the controller to deliberately choose grasping points depending on the object shape. For example, a rectangular shaped object needs grasping points different from a triangular shaped one. For this experiment, we used the objects of the RGB stacking benchmark (Lee et al., 2021), see figure 7, right. Please note, that the distribution of shapes is different for different colors. This explains the varying success rates for different colors, e.g. the green objects are more challenging to lift in average.
@@ 4.5 Real Robot Experiments - Teacher/ Student setup
The dispatcher in this case is restricted to find colored objects and advise the executor accordingly (by setting the right mask), but it is obvious how to extend this to more general objects, e.g. exploiting the recent advancement in general segmentation algorithms, e.g. (Oquab et al., 2023).
As a teacher policy, we took a policy that has been fully self-learned on real robot experience to stack the red object on the blue object on the five different object sets of the RGB stacking benchmark (Lampe et al., 2023). Note that this was learned from ground off directly from pure experience and no additional simulation data or tele-operation data was used. The performance of the original policy was about 96% in average over all ’red on blue’ object sets 1-5.
@@ 4.6 The power of the dispatcher
Triple stacking means stacking three objects on top of each other. This has been tried by other efforts using a single policy (Bousmalis et al., 2023). Here, we modify the dispatcher and ask to stack object 1 on object 2, followed by stack object 3 on object 1. Again, the executor is the basic D/E stacking policy. The observed success rate is $38\%$. The main reason for failure here is that the second stack is not high enough (this was not seen during the original training phase). However, the initial performance is strong, and finetuning the executor with accordingly collected data will solve this task.

### v2

@@ Abstract
Humans instinctively neglect irrelevant details when solving complex decision-making problems in environments with unforeseeable variations. This abstraction process appears to be a vital property for biological systems, helping to abstract away” unnecessary details and boost generalization. In this work, we introduce the Dispatcher/Executor (D/E) principle for the design of multi-task Reinforcement Learning controllers. It proposes partitioning the controller into two entities: one that understands the task (the Dispatcher) and one that computes the controls for the specific device (the Executor) - connected by a strongly regularized communication channel. The core rationale behind this position paper is that changes in structure and design principles can improve generalization properties and drastically enhance data efficiency. It is, in a sense, a “yes, and…” response to the current trend of using large neural networks trained on vast amounts of data to rely on emergent generalization properties. While we agree on the power of scaling - in the sense of Sutton’s “bitter lesson” - we provide evidence that considering structure and design principles remains a valuable and critical component, particularly when data is not abundant and infinite, but a precious resource. A video showing the results can be found at https://sites.google.com/view/dispatcher-executor.
@@ 1 Introduction
Reinforcement Learning (RL) has evolved from single-task settings to multi-task scenarios, where a single agent must solve one of many potential tasks. Such requirements are ubiquitous in robotics, where a manipulator must handle diverse objects or a locomotion platform must navigate to varying goals. Multi-task challenges also arise in classical control, such as shaping fusion plasma (degraveNature2022) or general set-point control (hafnerriedmiller2011). The prevailing approach typically relies on a monolithic neural network, conditioned on a task specification to modulate its behavior.
However, effective multi-task control demands two distinct types of knowledge: semantic task understanding (e.g., identifying which objects are relevant or what manipulation is required) and mechanical execution (e.g., specific knowledge of the device’s kinematics and dynamics). Standard monolithic architectures conflate these requirements, forcing a single network to simultaneously learn general world knowledge and specific actuation skills.
In this work, we argue that multi-task RL benefits significantly from a bi-partite architecture that explicitly separates these concerns: a Dispatcher that parses the semantic intent, and an Executor that generates the control actions. Crucially, the communication interface between these modules should enforce a strict information bottleneck, filtering out task-irrelevant details. This forced abstraction allows the same Executor to be reused across a broad diversity of visual scenes, thereby boosting generalization—validating the principle that “less is more.”
We formalize this Dispatcher/Executor (D/E) principle and present a concrete implementation within the domain of robotic manipulation. We provide extensive empirical evaluations demonstrating significant benefits across a range of tasks in both simulation and on a real robot. Notably, the D/E principle can also be applied retrospectively; we demonstrate this “hindsight transfer” by decomposing an existing real-world policy (trained for stacking specific colored blocks) into a D/E structure capable of general object stacking.
@@
  • Introduces the Dispatcher/Executor (D/E) principle as a design paradigm for scalable multi-task RL controllers.
  • Proposes a concrete architecture for D/E in robotic manipulation, serving as both a proof of principle and a practical implementation guide.
  • Provides empirical evaluations in simulation and on real hardware, demonstrating substantial gains in data efficiency and generalization for multi-task RL.
@@ 2 State of the Art
The core claim of Rich Sutton’s insightful ’Bitter Lesson’ paper (suttonBitter2019) is simple and compelling: ’general methods that leverage computation are ultimately the most effective’. This has been demonstrated several times in the history of AI, e.g. in game playing (silver2016mastering), computer vision (KrizhevskySH17), or more recently, in large language models (VaswaniSPUJGKP17): huge amounts of data trained in huge neural networks with huge amounts of compute lead to unprecedented and surprising generalisation (openai2023gpt4), while the underlying training principle is general and relatively simple.
@@
Given the big success in language, it is evident to investigate, if the same principles of scaling up to large amounts of data and the use of large networks show the same generalisation effects in control. The GATO (reed2022generalist) approach demonstrated that a reasonably large transformer style neural network is able to show good control performance with a single network on a variety of heterogeneous tasks from Atari to robotics. RoboCat (bousmalis2023robocat) used a related approach focusing on robotic control with the final goal to train a versatile model that can transfer to arbitrary new tasks. RT-1 (brohan2022rt) and RT-2 (zitkovich2023rt) start from a pre-trained language model and therefore have a built-in semantic understanding of tasks and the environment. All of the above models are trained on vast amounts of data using a transformer-style controller architecture with an additional input pathway to communicate the task. Since these methods are by default extremely data-hungry, data-augmentation is one path to generate even more data out of the data collected (yu2023scaling) and by this enforce generalisation within the distribution of the data generated.
@@
In ahn2022i it is shown that a pre-trained language model and a set of pre-trained skills can be efficiently combined by learning to sequence these skills, based on the assumption that language model and the skills share some common understanding of the main task. While this gives some form of general task execution in known environments, a large amount of data has to be used to make sure the pre-trained skills will be able to execute on a sufficient enough set of variations of the environment.
@@
Another scheme that is very common for data efficient multi task approaches is to learn some form of task embedding that allows to generalise in the task space (e.g. hausman2018learning; nachum2018dataefficient). While some of these ideas can be combined with the D/E approach, these methods solely focus on the command part in the communication channel. Therefore the expected generalisation by abstracting the observation and the command, as we demonstrate here, is missing.
@@
While we agree that scaling is one way to achieve more and more general controllers, we at the same time believe that even large scale systems should be ’data-efficient’, i.e. should make the best use of the data available. The core thesis of this paper is that incorporating structure into the architecture and the communication channel can drastically enhance the generalisation capabilities ’by design’. In particular, enforcing the right representation within the controller seems to be an important enabler. Recent examples of this line of research can be found in levine2016endtoend; pinneri2023equivariant; groth2021goal.
@@
The idea of a two (or more) level architectures for control has been explored in many different variations (e.g. vezzaniSkills2022; zhang2021hierarchical; riedmiller2018learning; hafner2022deep. The D/E principle intentionally defines the separation from the viewpoint of the type of knowledge required to run the components: a dispatcher component that understands the task through general world knowledge (that can be trained on general data like text, images or videos), and an executor component that specifically understands the device (trained through active interaction with the device). The restrictions on the communication channel enforce an abstract information exchange between these types of knowledge that boosts generalisation. We believe that in particular in this new era of large language models that are able to represent general world knowledge, this structural principle can play an important role to boost learning control on real robots, where data-efficiency with respect to device interaction is a key requirement.
@@ 3.3 Workflow and learning within the D/E architecture for manipulation
Learning in a D/E architecture is currently restricted to the executor. Learning is not different from the standard monolithical controller, apart from the fact, that the input is represented differently. The executor can therefore be trained in many different ways. In the following we investigate Reinforcement Learning from scratch as well as the distillation of a previously learned policy to a D/E architecture in a student/teacher setup. Another interesting direction, which is beyond the scope of this paper, is the investigation of learning multiple variants of representations in an off-line RL setting of pre-collected data (lange2012batch; riedmiller2021collect).
@@ 4 Empirical investigations of targeted properties
The dispatcher-executor architecture (see figure 1) uses a hardcoded dispatcher and a learned executor. The dispatcher gets the task specification (e.g. ’lift green object’ or ’stack red on blue object’) and then segments the according target objects - in this case using simple color segmentation (however, it is straightforward to replace this by more capable segmentation schemes, e.g. oquab2023dinov2. In addition, it filters the whole scene using an edge operator. Then, the executor is called, providing the respectively filtered images as separate arguments. If not otherwise noted, ’D/E’ refers to the default setting of using a segmentation mask for the objects, and the edge operator for the overall scene (as described in 3.2).
For training both the standard architecture and the executor in the D/E case, a distributional variant of MPO abdolmaleki2018maximum, with a Mixture-of-Gaussian critic was used. The agent receives the images of three cameras placed around the robot’s workspace, as well as proprioceptive observations. Both architectures employ a three-block ResNet that computes an eight-dimensional embedding vector for each pixel input. These embeddings are concatenated with all proprioceptive observations and fed through a three layer MLP for computing the agent’s five-dimensional actions.
@@ 4.3 Multi-task 2: Lifting one out of three objects with multiple shapes (simulation)
One critical question with a reduced representation is, whether the carried information is still rich enough to cope for desired variances in the task setup. To look a bit deeper into this, in the following the objects are additionally varied in shape. This problem is considerably more complicated than grasping cubes, forcing the controller to deliberately choose grasping points depending on the object shape. For example, a rectangular shaped object needs grasping points different from a triangular shaped one. For this experiment, we used the objects of the RGB stacking benchmark (lee2021beyond), see figure 7, right. Please note, that the distribution of shapes is different for different colors. This explains the varying success rates for different colors, e.g. the green objects are more challenging to lift in average.
@@ 4.5 Real Robot Experiments - Teacher/ Student setup
The dispatcher in this case is restricted to find colored objects and advise the executor accordingly (by setting the right mask), but it is obvious how to extend this to more general objects, e.g. exploiting the recent advancement in general segmentation algorithms, e.g. (oquab2023dinov2).
As a teacher policy, we took a policy that has been fully self-learned on real robot experience to stack the red object on the blue object on the five different object sets of the RGB stacking benchmark (lampe2023mastering). Note that this was learned from ground off directly from pure experience and no additional simulation data or tele-operation data was used. The performance of the original policy was about 96% in average over all ’red on blue’ object sets 1-5.
@@ 4.6 The power of the dispatcher
Triple stacking means stacking three objects on top of each other. This has been tried by other efforts using a single policy (bousmalis2023robocat). Here, we modify the dispatcher and ask to stack object 1 on object 2, followed by stack object 3 on object 1. Again, the executor is the basic D/E stacking policy. The observed success rate is $38\%$. The main reason for failure here is that the second stack is not high enough (this was not seen during the original training phase). However, the initial performance is strong, and finetuning the executor with accordingly collected data will solve this task.
4.7 Open-Vocabulary Semantic Dispatcher
Figure 10: Overview of the open-vocabulary Dispatcher-Executor architecture. The generic Dispatcher processes camera inputs through two parallel streams: a semantic stream using the pre-trained OWL-ViT to generate high-level segmentation masks based on natural language queries, and a structural stream using standard Sobel filters to compute ed ge maps. These task-agnostic outputs are fused into multi-channel augmented images serving as the exclusive input for the Executor. The Executor, implemented as a DMPO agent with generic CNN and MLP encoders, predicts manipulation actions relying solely on these structural outlines without perceiving raw RGB pixels.
In this series of experiments, we introduce a generic, open-vocabulary Dispatcher based on the pre-trained OWL-ViT Vision Transformer (10.1007/978-3-031-20080-9_42). This approach moves beyond task-specific, handcrafted detectors, allowing the generic Dispatcher to semantically interpret the scene based on natural language queries. The Dispatcher receives RGB inputs from available scene overview cameras and processes them through two parallel streams: a structural stream utilizing standard Sobel filters to compute generic, low-level edge maps, and a semantic stream where OWL-ViT, prompted by one or more natural language keywords, generates high-level segmentation masks for specified objects. These streams are fused into a strongly regularized communication channel represented as multi-channel augmented images for each view. The primary channel contains the raw, task-agnostic edge map, while subsequent channels contain this same edge map masked by the semantic segmentations corresponding to each generic text query. Consequently, the Executor—implemented as a DMPO agent with standard shallow generic encoders—never perceives raw RGB pixels; it relies solely on structural outlines highlighted by the Dispatcher according to the immediate semantic goal.
We validate this instantiation on a multi-object robotic manipulation task using the Aloha bi-manual platform. Four diverse fruits—a banana, pear, orange, and apple—are instantiated on the tabletop with randomized positions and orientations. The generic Dispatcher operates using only two fixed overview cameras (overhead and "worm’s-eye" generic views), without recourse to wrist-mounted sensors. The generic agent must identify, reach, and lift a specific target fruit from the cluttered table, with the target specified solely via a natural language keyword passed to the Dispatcher. The task utilizes a staged generic reward function that incentivizes reaching the target prop and subsequently lifting it above a predefined height threshold.
4.8 Experimental Evaluation: Semantic Decoupling and Geometric Transfer
To validate the hypothesis that the D/E architecture effectively compartmentalizes semantic knowledge from mechanical control, we conducted a series of zero-shot transfer experiments. We compare our Open-Vocabulary D/E agent against a strong baseline: a standard, monolithic Multi-Task RL agent (end-to-end RGB) trained with the same MPO algorithm and network capacity.
Our experiments utilize four target objects with distinct shapes and textures: an Apple (round, red), an Orange (round, orange), a Banana (elongated, yellow), and a Pear (irregular/conical, green). We focus on three specific transfer scenarios:
Table 1: Quantitative evaluation of zero-shot generalization. We compare the success rates of the Standard RGB baseline and the proposed D/E agent across three types of distribution shifts: visual background changes (Rows 1-2), visual clutter (Rows 3-4), and semantic transfer to novel target objects (Rows 5-12).
Exp Train Lift Task Test Lift Task Condition Method Success Rate
1 Apple Apple diff. background Zero-shot to different background Standard RGB 3.27%
2 Apple Apple diff. background Zero-shot to different background D/E 100%
3 Apple Apple + clutter Zero-shot with added clutter Standard RGB 7.62%
4 Apple Apple + clutter Zero-shot with added clutter D/E 98.71%
5 Apple Orange Zero-shot to similar shape diff. color Standard RGB 0%
6 Apple Orange Zero-shot to similar shape diff. color D/E 100%
7 Apple Pear Zero-shot to different shape D/E 91.67%
8 Apple Banana Zero-shot to very different shape D/E 16.24%
9 Banana Pear Zero-Shot to different shape D/E 87.94%
10 Apple $\lor$ Banana Pear Zero-Shot to different shape D/E 98.23%
11 Pear Banana Zero-shot to very different shape D/E 23.35%
12 Pear $\lor$ Apple Banana Zero-shot to very different shape D/E 40.02%
1. Invariance to Visual Background Shift: We evaluated the ability to ignore irrelevant visual noise. We train both the D/E agent and the Baseline agents on the “Apple lift” task. Both agents are trained in an environment having an “Office” background (youtu.be/8ZskIr7zurA). We then evaluated them zero-shot against a uniform dark blue background, keeping the robot and target object identical (youtu.be/mzco_zMMVSc). The results (Table 1, Rows 1-2) demonstrate the standard RGB agent’s inability to generalize under distribution shifts induced by background changes. The agent overfits to spurious correlations, illustrating a common failure mode of standard end-to-end learning. The Baseline agent suffers a performance collapse as the success rate drops to 3.27%. In contrast, the D/E principle mandates a regularized communication channel that allows the performance to be unaltered by background noise. In the presented instantiation, the OWL-ViT based Dispatcher explicitly filters out task-irrelevant background information. This structural constraint ensures the Executor perceives the uniform background as functionally identical to the complex office scene, preserving high performance without the need for domain randomization or retraining.
2. Robustness to Semantic Distractors: We evaluate the ability to ignore added clutter. We take both agents trained on the “Apple lift” task and apply them without any additional training on a test environment where four fruits are placed in the workspace simultaneously (youtu.be/ClcoUzsIR0M). The D/E agent succeeds in the vast majority of cases (Table 1, Row 4). The Dispatcher handles the visual complexity by silencing distractor objects. This ensures the Executor perceives a state functionally similar to a single-object scene, even if the target is partly occluded. Consequently, performance remains high, whereas the success rate of the standard end-to-end baseline deteriorates significantly (Table 1, Row 3).
3. Zero-Shot Generalization to Novel Objects: We evaluate the ability to transfer learned manipulation behaviors to semantically similar contexts. Specifically, target objects with similar geometry but different visual appearance (e.g., color and shape). We apply both agents, originally trained on the “Apple lift” task, to an “Orange lift” task without any additional training (youtu.be/OlH8PA9Q6IU). For the D/E agent, adaptation requires only updating the text prompt sent to the Dispatcher from “Apple” to “Orange”. In contrast, the Baseline agent, having overfitted to the specific texture and color of the apple, fails completely, achieving a 0% success rate on the orange (Table 1, Row 5). The D/E Executor, however, receives regularized input; since the semantic mask of the orange is geometrically similar to that of the apple, the Dispatcher treats it as a known object. This results in a 100% success rate without fine-tuning (Table 1, Row 6).
We extend this evaluation by performing additional zero-shot tests on target objects with increasing geometric dissimilarity from the target object used during training (Table 1, Rows 7-8). The D/E method, achieves a 91.67% zero-shot success rate on the “Pear lift” task, despite the pear’s distinct color and only partial geometric similarity to the apple (youtu.be/cu__ZcXYi4k). The “Banana lift” task achieves a 16.24% success rate (youtu.be/k2k2lqif_TE), even though the banana shares neither texture nor geometric similarity with the original target object. As expected, zero-shot performance correlates with the degree of geometric similarity to the original training object. Crucially, the D/E approach enables the retention of some success capabilities even when facing significantly different target objects.
Figure 11: Training efficiency comparison on the “Apple lift” task. The plot reports the success rate (mean $\pm$ standard deviation) as a function of the number of training episodes. The agent trained jointly with the “Pear lift” task (green) converges significantly faster than the agent trained on “Apple lift” in isolation (orange), validating the data efficiency achieved by the shared Executor.
Figure 12: Sequential transfer efficiency on the “Apple lift” task. We compare training from random initialization (orange) against fine-tuning from a policy pre-trained on “Banana lift” (blue) and a policy pre-trained on both “Banana lift” and “Pear lift” (red). Pre-training on a single task reduces sample complexity by over 50%, while pre-training on two tasks yields immediate 100% zero-shot success.
4. Generalization and Efficiency via Multi-Task Learning
We aim to demonstrate empirically that the Dispatcher and Executor modules acquire generic, reusable knowledge. Specifically, we hypothesize that an Executor shared across a higher number of tasks performed by the same embodiment acquires a more general representation of procedural (“how”) knowledge. We measure this generality through two metrics: training efficiency and zero-shot performance on unseen scenarios.
Training Efficiency. First, we train a shared Executor (from random initialization) jointly on two distinct tasks: “Apple lift” and “Banana lift”. While the Executor is shared, each task utilizes a specific OWL-ViT Dispatcher conditioned on a distinct text query (“Apple” or “Banana”). We limit the joint training to the same total computational budget used for the single-task “Apple lift” baseline. Consequently, although each task in the joint setup effectively sees only half the training samples, both achieve a 100% success rate. In contrast, training the “Apple lift” policy in isolation with this reduced sample size fails to converge (Figure 12). This confirms that a shared Executor improves data efficiency by acting as a repository for embodiment-specific knowledge that is transferred across tasks.
Zero-Shot Generalization. We further test the hypothesis that sharing the Executor improves zero-shot generalization to unseen objects. We compare three policies trained from random initialization with identical compute budgets: 1) Trained solely on “Apple lift”; 2) Trained solely on “Banana lift”; 3) Trained jointly on “Apple lift” and Banana “lift”, forcing the Executor to learn grasping primitives for both round and elongated geometries. We evaluate these policies zero-shot on the “Pear lift” task. As expected, the multi-task Executor achieves superior performance (Table 1, Rows 7, 9, 10). Transferring from the single-task “Apple” and “Banana” policies yields success rates of 91.67% and 87.94%, respectively. In contrast, the jointly trained Executor achieves 98.23%, reducing the error rate by more than a factor of four. To validate the generality of this phenomenon, we repeat the experiment targeting the banana for zero-shot evaluation, using policies trained on apples and pears (Table 1, Rows 8, 11, 12). While single-task transfer achieves only 16.24% and 23.35%, the jointly trained Executor reaches 40.02%, providing further empirical support for the proposed hypothesis.
Sequential Transfer Efficiency. Finally, we investigate the benefits of the shared Executor in a sequential transfer setting, distinguishing between learning from scratch and fine-tuning from pre-existing knowledge. We compare three training regimes for the “Apple lift” task: 1) training from random initialization, 2) fine-tuning from a policy pre-trained on the “Banana lift” task, and 3) fine-tuning from a policy pre-trained on both “Banana lift” and “Pear lift” tasks jointly. The results are illustrated in Figure 12. The baseline policy (random initialization) requires approximately 550,000 episodes to reach convergence. In contrast, fine-tuning from a single related task (“Banana lift”) dramatically accelerates learning, converging in less than half the number of samples. Most notably, the policy pre-trained on two tasks (“Banana” and “Pear”) demonstrates perfect immediate generalization, achieving a 100% success rate from the start of training. This progression—from slow learning, to rapid adaptation, to immediate zero-shot execution—provides compelling evidence that the Executor accumulates robust, reusable procedural knowledge as it is exposed to a wider variety of tasks.
Collectively, these experiments validate the core premise of the Dispatcher/Executor principle: that effective multi-task control relies on the clean separation of semantic intent from mechanical execution. By decoupling the “what” (Dispatcher) from the “how” (Executor), we break the dependency between perceiving a new context and learning the actuation primitives to act within it. As demonstrated, this architecture not only ensures robustness against visual distractors and background shifts but also allows the agent to apply existing control primitives to novel scenarios immediately upon identification. Consequently, the D/E framework effectively reduces the marginal sample complexity of learning new tasks, offering a scalable path toward general-purpose robotic behavior learning.